# A 4 - 8 GHz Kinetic Inductance Travelling-Wave Parametric Amplifier Using Four-Wave Mixing with Near Quantum-Limit Noise Performance

### 19 Feb 2024 v1

@@ Appendix B Noise Theory
Where $N_{in}(T)=\frac{1}{2}\coth\Big{(}\frac{\hbar\omega}{K_{B}T}\Big{)}$ is the input noise of the 50 $\mathrm{\Omega}$ terminations. $N_{qm}$ is the quantum noise and for $k_{B}T\ll\hbar\omega$ is equal to one half. $G_{PA},N_{A}$ and $G_{HEMT},N_{HEMT}$ are the gain and added noise of the KI-TWPA and the HEMT amplifier, respectively. $G_{w}$ is the additional effective warm amplification, and $N_{tot}$ is the total noise measured on the spectrum analyzer (SA).
@@
Improve it In Fig. 5.b, you can see the transmission ($S_{21}$) measurements of the cable in two scenarios. The first one is the bypassing of KI-TWPA, represented by the orange line, and the second one is the transmission through DPX2, KI-TWPA, and DPX3, represented by the blue line. We have calculated the loss between switch B1 and the KI-TWPA by subtracting the unpumped measurement from the bypass measurement and dividing it by two since we used identical diplexers with similar insertion loss. This calculated loss is plotted as a green line in Fig. 5.b as a function of frequency, and we have used it as L1 and L2 loss factors in the main text.

### 13 Mar 2024 v3

@@ Appendix B Noise Theory
Where $N_{in}(T)=(1/2)\coth\left(\hbar\omega/2k_{B}T\right)$ is the input noise from the 50 $\mathrm{\Omega}$ terminations. $N_{qm}$ is the quantum noise and for $k_{B}T\ll\hbar\omega$ is equal to one half. $G_{PA},N_{A}$ and $G_{HEMT},N_{HEMT}$ are the gain and added noise of the KI-TWPA and the HEMT amplifier, respectively. $G_{w}$ is the additional effective warm amplification, and $N_{tot}$ is the total noise measured on the spectrum analyzer (SA).
@@
In Fig. 5.b, you can see the transmission ($S_{21}$) measurements of the cable in two scenarios. The first one is the bypassing of KI-TWPA, represented by the orange line, and the second one is the transmission through DPX2, KI-TWPA, and DPX3, represented by the blue line. We have calculated the loss between switch B1 and the KI-TWPA by subtracting the unpumped measurement from the bypass measurement and dividing it by two since we used identical diplexers with similar insertion loss. This calculated loss is plotted as a green line in Fig. 5.b as a function of frequency, and we have used it as L1 and L2 loss factors in the main text.